OpenStudy (anonymous):

a=-2 , b=-1 , c=4 , d=3 Solve) (Square Root Of) a = ? Solve) bc

OpenStudy (anonymous):

\[\sqrt{-2} = \sqrt{2}i\] \[bc = b*c = (-1)(4) = -4\]